# Knowledge-Driven Feature Selection and Engineering for Genotype Data with Large Language Models

### v1

@@ Abstract
Predicting phenotypes with complex genetic bases based on a small, interpretable set of variant features remains a challenging task. Conventionally, data-driven approaches are utilized for this task, yet the high dimensional nature of genotype data makes the analysis and prediction difficult. Motivated by the extensive knowledge encoded in pre-trained LLMs and their success in processing complex biomedical concepts, we set to examine the ability of LLMs in feature selection and engineering for tabular genotype data, with a novel knowledge-driven framework. We develop FreeForm, Free-flow Reasoning and Ensembling for Enhanced Feature Output and Robust Modeling, designed with chain-of-thought and ensembling principles, to select and engineer features with the intrinsic knowledge of LLMs. Evaluated on two distinct genotype-phenotype datasets, genetic ancestry and hereditary hearing loss, we find this framework outperforms several data-driven methods, particularly on low-shot regimes. FreeForm is available as open-source framework at GitHub: https://github.com/PennShenLab/FREEFORM.
@@ Introduction
Predicting observable phenotypes from genotypic data has proven to be a monumental task in the field of genetics with diverse applications, ranging from personalized medicine 1 to genomic selection of crops2. The prediction methods attempt to harness vast amounts of genetic variants as features, such as single nucleotide polymorphisms (SNPs) detected from high-throughput sequencing data, to make predictions on the phenotype. Linear models that deal with the sparsity of genomics data with regularization have been developed to much success in the case of polygenic risk scores 1; 3. More sophisticated machine learning models have also been developed to capture the entangled epistatic relationships between genes and to predict complex traits 4.
@@
The difficulty, however, of modeling genotype data is substantial: (1) First, the topic of interest is not merely predictive (e.g. discovering causal variants5) and machine learning methods are known to identify spurious features as significant due to multicollinearity 6. Analyzing interaction terms—like epistatic relationships between variants—also introduces challenges such as a lack of interpretability when using complex models or multiple testing when examining higher-order interactions 7. (2) Furthermore, a table of genotype data could contain thousands or even millions of columns. This ‘curse of dimensionality’ can cause severe overfitting and amplify many issues, including those we’ve named above8. (3) Lastly, data is often limited in real genotype datasets, exacerbating existing concerns with overfitting.
@@
Feature selection and feature engineering (i.e. feature construction)9 can be crucial steps to mitigate these concerns. Data-driven methods, such as Lasso regression, have demonstrated great success in reducing the dimensionality of genotype datasets 10. Feature engineering has shown to improve predictive performance without using less interpretable models and help uncover interactions within the data 11; 12. However, these methods have their own issues: data-driven methods can struggle with small sample sizes and feature engineering is a laborious process that requires expertise to avoid multiple testing.
Recent advances in large language models (LLMs) 13; 14 have shown promise in addressing these challenges. LLMs’ extensive knowledge equipped with well-designed prompting strategies 15; 16; 17 has led to promising few-shot or zero-shot performance across various tasks 18; 19. Moreover, several studies have proposed to employ such few-shot or zero-shot capabilities of LLMs to perform feature preprocessing. Among them, Choi et al. 20 adopts extensive prior knowledge in LLMs to conduct feature selection and casual discovery. Jeong et al. 21 introduces three different feature selection pipelines with LLMs, all of which directly utilize the generated text output to rank the features’ importance. Some studies have explored more advanced pipelines. Liu et al. 22 introduces in-context evolutionary search (ICE-SEARCH), which involves recurrently updating the selected features via feature filtering of LLMs based on test scores in each feature set. Han et al. 23 employs LLMs to do feature engineering, generating conditional rules for each class label to train linear models.
@@
Observing the promising performance of LLMs, there has been a growing exploration of LLMs’ usage in biomedical applications, including genetics 24; 25; 26. Despite these innovations, most genetic studies have focused on gene-level data, likely due to the limitations of earlier LLMs. To our best knowledge, no studies have explored LLMs targeting variant-level data, except one that used API calls on NCBI databases to retrieve SNP information 27.
@@
Inspired by these recent explorations 21; 23; 28; 29; 30, we propose to leverage LLMs’ knowledge (intrinsic and augmented) 31; 32 and reasoning capabilities 15; 33 to select the most informative genetic variants and generate novel features. We develop a knowledge-driven framework FreeForm, Free-flow Reasoning and Ensembling for Enhanced Feature Output and Robust Modeling, designed from ground principles in ensembling16 and “free-flow reasoning”34 to best leverage the expertise of LLMs. In the LLM-enabled framework, we implement scalable feature selection strategies that can process a large number of variants and feature engineering approaches that focus on interaction terms which are more interpretable. We evaluate FreeForm on two real genotype datasets, genetic ancestry and hereditary hearing loss, and compare with data-driven and LLM-enabled methods. In particular, we focus on few-shot settings where traditional data-driven methods struggle due to limited sample size but LLMs have shown surprising generalizability 35; 36. In our study, we expect challenges when applying LLMs to tabular genotype data due to their limited semantics —column names are variant IDs and values indicate the number of minor allele copies (0,1,2). Thus, we analyze the effect of retrieval augmentation and domain-specific serialization when prompting.
@@
Our results highlight FreeForm’s potential to address three challenges present in modelling genotype data: (1) Our method enhances predictive modelling while maintaining interpretability, (2) reduces the dimensionality of the dataset and (3) excels in few-shot regimes compared to data-driven methods by grounding our methods in knowledge.
@@ Methods
Our goal for feature selection is to find a subset of variants $S^{\prime}\subseteq S$ of size $d^{\prime}\ll d$ through pre-trained LLMs such that a downstream model $f$ trained on $D_{S^{\prime}}=\{(\mathbf{x}^{i}_{S^{\prime}},y^{i})\}^{N}_{i=1}$ can make efficient predictions of the phenotype with this smaller subset of variants. The LLM, denoted as a stochastic operator $\mathcal{T}$ parameterized by $\theta$ and given prompt $p$, can be formally described as:
@@
where $\theta$ captures the stochasticity inherent in the LLM. The proposed selection method is model agnostic unlike model-based rankings; in the downstream analysis, we use two classic models, Random Forest (RF) and Logistic Regression (LR), for demonstration. Also, previous LLM-based feature selection methods have not been tested in the setting of variants-level data where hallucination can be more likely 37 and high dimensionality is an issue. While advancements in context window size have mitigated some issues, new challenges have emerged 38; 39; to address these challenges, we design our feature selection strategies to scale for high dimensions while remaining token-efficient.
@@
We ask the LLM to determine whether each of the $d$ variants is relevant to the task, requesting a simple “Yes” or “No”, yielding a subset of variants. This set may still be large, so we adjust the language of the prompt appropriately (e.g. “potentially relevant”) based on how many are filtered.
@@
is the first strategy we employ for selecting $d^{\prime}$ features from the filtered set of variants. We begin by randomly partitioning the variants into buckets of approximately equal size, each containing around 50 to 100 variants (where this number is a hyper-parameter). This is to avoid issues of losing information when contexts become too large38. Then, each bucket is independently passed to the LLM, tasked with selecting the $d^{\prime}$ most relevant variants; we select $d^{\prime}$ at every step in the case that the relevant features are concentrated in a single bucket. The $d^{\prime}$ selected variants from each bucket are then merged together and the process repeats.
We observe that the selection process is sensitive to the order in which variants are presented to the LLM. Thus, for a particular bucket, we conduct multiple iterations, randomizing the order of variants in each iteration, and ultimately select much more than $d^{\prime}$ features. This approach naturally integrates self-consistency, retaining the top $d^{\prime}$ variants that appear the most frequently across iterations. To enhance diversity in feature selection, we introduce a non-zero temperature setting of 0.3. This process is repeated until we have $d^{\prime}$ features from a single bucket. To enhance the LLM’s precision and robustness during the crucial, final selection, we enable the LLM with chain-of-thought (CoT) prompting, increase the temperature to 0.7, and increase the number of iterations to ten.
@@
is the second strategy we employ. Starting with the filtered set of $f$ variants, we task the LLM with identifying the single most relevant variant with CoT. After each selection, the chosen variant is removed from the pool, and the process is repeated. Initially, we perform this extraction without any ensembling, performing the extraction once per iteration. We find that the top few features are easy for the LLM to identify, but the task becomes more challenging as it becomes ambiguous which variants are more significant. As the selection becomes more challenging, we apply self-consistency: extracting a variant several times, in an increasing manner as we near the end of the selection, per iteration with a non-zero temperature. As we incorporate self-consistency and repeat steps, we find that this also mitigates the LLM’s tendency to return the features that were already selected. In cases where the LLM still fails to identify a valid variant, which occurs frequently towards the end, we implement an exception handling mechanism that retries the selection process.
@@
Next, the goal for feature engineering is to transform the dataset $D_{S^{\prime}}$ into $K$ transformed datasets $D_{k}=\{(\mathbf{x}^{i}_{S^{\prime}_{k}},y^{i})\}^{N}_{i=1}$, where $S^{\prime}_{k}\supseteq S^{\prime}$. To create these transformed datasets, we use the LLM $\mathcal{T}$ with prompt $p^{\prime}$ to engineer new sets of features, which we then augment onto $D_{S^{\prime}}$. In the instructions of $p^{\prime}$, we include a subset of examples, $R=\{(\mathbf{x}^{i}_{S^{\prime}},y^{i})\}_{i\in R\subseteq D_{S^{\prime}}}$, to enhance the LLM’s prior knowledge with in-context learning 35.
Tabular data, however, may not be in a format that the LLM easily understands. We construct a function $\mathtt{Serialize}$ that takes these IDs and converts each row in the table to a textual description like “The $s_{1}$ variant of the person has $x^{i}_{1}$ minor alleles... The $s_{d^{\prime}}$ variant of the person has $x^{i}_{d^{\prime}}$ minor alleles.” with a phenotype class string “$y^{i}$”. Since the performance of LLMs is sensitive to the text input, we examine different serialization templates and prompts. Furthermore, since some of the variant IDs may not be known to the LLM, we also explore using retrieval-augmentation  40 to add information from external sources to enhance the pre-trained LLM.
The LLM, $\mathcal{T}$, generates new features that are added back onto $S^{\prime}$ to give $S^{\prime}_{k}$:
$S^{\prime}_{k}=S^{\prime}\cup\mathcal{T}\left(p^{\prime};\theta\right),\text{where }p^{\prime}\text{ includes }\{\mathtt{Serialize}(\mathbf{x}^{i}_{S^{\prime}},y^{i},S^{\prime})\}_{i\in R}$ (2)
After generating these datasets, we train a set of models =$\mathcal{F}=\{f_{k}\}^{K}_{k=1}$, where each $f_{k}\in\mathcal{F}$ is trained on a specific $D_{k}$, potentially capturing different hypotheses about how the variants relate to the phenotype. During inference, we average the class probabilities $\mathbf{p}(f_{k}(\mathbf{x}))$ predicted by each model $f_{k}$ for a particular input $\mathbf{x}$. The final prediction $\hat{y}$ is made by selecting the class $i$ with the highest averaged probability:
@@
Our aim is to leverage the diverse feature representations generated by the LLM, mitigating the risks associated with overfitting in low-shot regimes and thereby improving its generalization to unseen data. To prevent the loss of interpretability caused by complex methods, we also limit our engineering to traditional interaction terms in our instructions of $p^{\prime}$. We now discuss some of the key steps in this process in depth.
@@
Traditionally, LLMs are guided by prompts such that its output is structured, enabling it to be parsed through methods such as regular expressions. However, we deliberately allow the LLM to produce unstructured output, recognizing that enforcing a rigid structure can diminish the depth and quality of the LLM’s CoT reasoning 34. To guide the LLM, we provide an example of a “correctly” engineered feature 41, which we self-generate with an LLM and then manually verify, stabilizing the output.
@@
To mitigate overfitting, we repeat the feature engineering process $K$ times to generate all $D_{k}$ which we’ll use to train $K$ classifiers. Similar to FeatLLM 23, we incorporate additional ensembling and order bias mitigation strategies such as bagging and order shuffling, where we pass a random subset of $|R|\leq N$ samples to the LLM; we limit the number of samples $|R|$ to $16$. We do this to further diversify the LLM output for each iteration, and simultaneously avoid exorbitant usage of the context window. Our free-flow reasoning approach further contributes to the diversity of the output, as we naturally allow the LLM to determine the number of features that will be added to $S^{\prime}$ in its unstructured generation. By setting a non-zero temperature of 1, we ensure that each iteration further produces a varied set of features, especially in type. The resulting $k$ transformed datasets form the basis for the ensembled model.
Notably, our method is model-agnostic, de-coupled from the classifier. For demonstration purposes, we apply it to two widely used models in our downstream evaluation: Random Forest and Logistic Regression.
Results and Discussion
@@
Our experiments involve two datasets: the Genomic Ancestry Dataset42 and the Hereditary Hearing Loss Dataset43. The Genomic Ancestry Dataset is derived from the 1000 Genomes Project (1KGP). We focus on determining the superpopulation ancestry phenotype (African, American, East Asian, European, and South Asian). In particular, we used a curated set of 10,000 SNPs predefined by GRAF to pinpoint ancestry markers 44, and we’ll discuss how we addressed the issues that arise from their quality control (QC). After QC and preprocessing, the dataset includes 2,403 subjects and 8,688 variants as columns in the rsID format, a standard identifier used by dbSNP.
The Hereditary Hearing Loss Dataset is considerably smaller, comprising of 1,209 samples and 144 variants as columns, employing a nomenclature system used by ClinVar. This dataset is notably imbalanced, with approximately 75.9% of the samples classified as “Yes” (indicating the presence of hereditary hearing loss) and 24.1% as “No.” To our best knowledge, these two are the only open-access genotype datasets available online.
@@
For feature selection, we conduct comparisons with four baseline methods. The first three baselines are conventional machine learning approaches: (1) LASSO (Least Absolute Shrinkage and Selection Operator), (2) PCA (Principal Component Analysis) and (3) RF-based Gini Importance where we fit a Random Forest on the training data and rank the features by their Gini importance. In our case, the specification of a fixed number of features is required for our study. LASSO, however, reduces an arbitrary number of coefficients to zero during model training. Similarly, PCA provides a weighted sum of features, rather than selecting individual features. To adapt these methods to our requirements, we optimized LASSO over a grid of hyperparameters and selected the $d^{\prime}$ features with the largest coefficients in the highest-performing model. For PCA, we consider the common approach of taking the top loadings of the first principle component. Lastly, we include (3) LLM-Select21 (we use their LLM-RANK prompts).
For feature engineering, we compared our approach against five baselines. The first three are traditional machine learning methods: (1) Logistic Regression, (2) Random Forest and (3) XGBoost. Additionally, we includ more recent baselines, (4) TabPFN 45, a Transformer trained to do supervised classification for small tabular datasets, and (5) FeatLLM23 which also leverages LLMs to do feature engineering but limits themselves to conditional rules (e.g. variant 1 is 0) for a linear model.
@@
Our FreeForm framework utilizes GPT-4o (2024-05-13) as the primary LLM backbone, particularly for tasks requiring advanced reasoning capabilities, such as automating feature engineering and selecting relevant features. For more routine tasks, including parsing output and writing Python functions, we employ GPT-3.5-turbo with a temperature of 00 which offers a cost-effective solution that meets the performance requirements for these functions.
We employ an ensemble of $K=20$ models, striking a balance between cost-effectiveness and model performance, noting that performance gains diminish beyond this point. In replicating baselines, machine learning models were implemented using Python’s scikit-learn library. Hyperparameters were optimized using grid search combined with $k$-fold cross-validation, with $k$ set to either 2 or 4, ensuring that the training set includes at least one example of each class. For other methods, such as FeatLLM and TabPFN, we used default parameters with slight adjustments for fair comparison (e.g., using 15 conditions instead of 10 for FeatLLM). Also, for the evaluation of all feature engineering methods, we choose one of the feature sets generated by hierarchical selection.
@@
Figure 3: Evaluation of Feature Selection Methods with Different Downstream Classifiers
Our findings indicate that LLM-driven methods significantly outperform data-driven approaches for feature selection, in low-shot regimes. Notably, in the genomic ancestry task, our framework achieves performance comparable to LASSO with just 10 shots, whereas LASSO requires 80 shots to achieve similar results. In hearing loss, we observe that the performance gap is smaller and our advantage remains until 16 shots, likely due to the dataset’s much lower dimensionality. This exemplifies the strength of our methods in few-shot regimes such as genotype data.
For feature engineering, we conduct our experiments five times with cross-validation, limiting to few-shot settings where $N<120$. In Fig 4, FreeForm consistently ranks at the top or matches the performance of baseline models. In the genomic ancestry task, our framework enhances the performance of both Logistic Regression and Random Forest, especially in the lower-shot scenarios. Additionally, our approach outperforms recent models such as FeatLLM and TabPFN. Notably, as the number of shots approaches 80, the performance gap between our methods and the baselines decreases. In the hearing loss task, our framework notably enhances the performance of Logistic Regression on higher-shot scenarios. While Random Forest does not benefit from the engineered features, our approach remains competitive. This outcome suggests limited effectiveness of the interaction terms we engineer for complex models like Random Forests.
Figure 4: Evaluation of Feature Engineering on the Two Datasets

### v2

@@ Abstract
Predicting phenotypes with complex genetic bases based on a small, interpretable set of variant features remains a challenging task. Conventionally, data-driven approaches are used for this task, yet the high dimensional nature of genotype data makes the analysis and prediction difficult. Motivated by the biomedical knowledge encoded in pre-trained LLMs and the emerging applications for genetics, we set to examine the ability of LLMs in feature selection and engineering for tabular genotype data, with a novel knowledge-driven framework. We develop FreeForm, Free-flow Reasoning and Ensembling for Enhanced Feature Output and Robust Modeling, designed with chain-of-thought and ensembling principles, to select and engineer features with the intrinsic knowledge of LLMs. Evaluated on two distinct genotype-phenotype datasets, genetic ancestry and hereditary hearing loss, we find this framework outperforms several data-driven methods, particularly on low-data regimes. FreeForm is available as open-source framework at GitHub: https://github.com/PennShenLab/FREEFORM.
@@ Introduction
Predicting observable phenotypes from genotype data has proven to be a monumental task in the field of genetics, with diverse applications ranging from personalized medicine 1 to genomic selection of crops2. Vast amounts of genetic variants, such as single nucleotide polymorphisms (SNPs) from high-throughput sequencing data, are often harnessed to make predictions on the phenotype. To deal with the sparsity of this data, linear models with regularization have been developed to much success in the case of polygenic risk scores 1; 3. Furthermore, machine learning models have been developed to capture the entangled epistatic relationships between genes and to predict complex traits 4.
@@
The difficulty, however, of modeling genotype data is substantial: (1) First, the topic of interest is not merely predictive (e.g. discovering causal variants5) and machine learning methods are known to identify spurious features as significant due to multicollinearity 6. Analyzing interaction terms—like epistatic relationships between variants—also introduces challenges such as a lack of interpretability when using complex models or multiple testing when examining higher-order interactions 7. (2) Furthermore, a table of genotype data could contain thousands or even millions of columns. This ‘curse of dimensionality’ can cause severe overfitting and amplify many issues, including those we’ve named above8. (3) Lastly, data can be limited in real settings, exacerbating existing concerns with overfitting.
@@
Feature selection and feature engineering (i.e. feature construction) can be crucial steps to mitigate these concerns. Data-driven methods, such as Lasso regression, have demonstrated great success in selecting features 9. Feature engineering improves predictive performance without resorting to complex models and can help uncover interactions between features 10. However, these methods have their own issues: data-driven methods can struggle with small sample sizes and feature engineering is a laborious process that requires expertise to avoid multiple testing.
Recent advances in large language models (LLMs) have shown promise in addressing these challenges. With their remarkable performance across various tasks, LLMs have established themselves as powerful tools across many domains 11; 12. A key strength of LLMs lies in the knowledge they acquire through pre-training, enabling them to act as domain experts; recent LLMs have showed extensive understanding of biomedical concepts 13; 14; 15. Furthermore, LLMs can be greatly enhanced with well-designed prompting strategies 16; 17; 18. Chain-of-thought prompting (CoT)  16 improves the reasoning of LLMs by encouraging step-by-step problem solving. Self-consistency 17 addresses the naive greedy decoding used in CoT prompting by selecting the most consistent outcome across multiple reasoning paths.
Several studies have proposed to employ such capabilities of LLMs to perform feature preprocessing. Among them, Choi et al. 19 adopts the notion of prior knowledge in LLMs to conduct feature selection and casual discovery. Jeong et al. 20 examines three types of selection strategies with LLMs e.g. ranking vs. scoring features. Hollmann et al. 21 has an agentic approach, using the LLM to generate Python code that creates features in an iterative fashion based on cross-validation feedback. Han et al. 22 employs LLMs to do feature engineering, generating conditional rules for each class label (e.g. age $>$ 21 increases the logits for label 00) and repeating this several times to form an ensemble.
@@
Due to the promising performance of LLMs, there has also been a growing exploration of LLMs’ usage in biomedical applications, especially genetics 23; 15; 24; 25. Despite these innovations, most genetic studies have focused on gene-level data, likely due to the limitations of earlier LLMs. To our knowledge, no prior work has explored LLMs on variant-level data, except one that used API calls on NCBI databases to retrieve SNP information 26. While this approach compensates for LLMs’ inability to pass the GeneTuring benchmark 27, we argue the benchmark is an inadequate test of their utility, as it evaluates gene-SNP association by randomly sampling 100 SNPs from hundreds of millions. This setup overlooks how well LLMs leverage their knowledge of known variants (Fig. 5), a gap we address in this study.
@@
Inspired by these recent explorations 20; 22; 23; 28, we propose to leverage LLMs’ knowledge, intrinsic and augmented, and reasoning capabilities 16; 29 to select the most informative genetic variants and generate novel features. We develop a knowledge-driven framework FreeForm, Free-flow Reasoning and Ensembling for Enhanced Feature Output and Robust Modeling, designed from ground principles in ensembling (self-consistency)17 and “free-flow reasoning”30 to best leverage the expertise of LLMs. In the LLM-enabled framework, we implement scalable feature selection strategies that can process a large number of variants and feature engineering approaches that focus on interaction terms which are more interpretable. We evaluate FreeForm on two real genotype datasets, genetic ancestry and hereditary hearing loss, and compare with data-driven and LLM-enabled methods. In particular, we focus on few-shot (i.e. few data samples) settings where data-driven methods struggle due to limited sample size but LLMs have shown surprising generalizability 31; 32. In our study, we expect challenges when applying LLMs to tabular genotype data due to their limited semantics—column names are variant IDs and values indicate the number of minor allele copies (0,1,2). Thus, we analyze the effect of retrieval augmentation and domain-specific serialization when prompting.
@@
Our results highlight FreeForm’s potential to address three challenges present in modelling genotype data: Our method (1) enhances prediction while upholding interpretability, (2) reduces the dimensionality of the dataset, and (3) excels in low-shot regimes compared to data-driven methods by grounding ours in knowledge. Furthermore, we challenge the notion that LLMs lack knowledge of genetic variants by novelly applying them in this domain.
@@ Methods
The goal of feature selection is to identify a subset $S^{\prime}\subseteq S$ with $d^{\prime}\ll d$ variants such that a downstream model $f$, trained on $D_{S^{\prime}}=\{(\mathbf{x}^{i}_{S^{\prime}},y^{i})\}^{N}_{i=1}$, can make efficient, interpretable predictions. We use an LLM, modelled as a stochastic operator $\mathcal{T}$ parameterized by $\theta$ and given prompt $p$, to output a selected subset of features $S^{\prime}$:
@@
Notably, the proposed selection method is model agnostic unlike model-based rankings. Furthermore, previous LLM-based feature selection methods have not been tested in the setting of variants-level data where hallucination can be more likely 33 and high dimensionality is an issue. While context windows have become longer, new challenges have emerged 34; 35 in which naive usage of the context would be ineffective; to address these challenges, we design our feature selection strategies to scale for high dimensions while remaining token-efficient. In the downstream analysis, we train two models, Random Forest and Logistic Regression, on the selected features to evaluate their quality.
@@
We first ask the LLM to determine whether each of the $d$ variants is relevant to the task, requesting a simple “Yes” or “No”, yielding a subset of variants. This set may still be large, so we adjust the language of the prompt appropriately (e.g. “potentially relevant”) based on how many are filtered.
@@
is the first strategy we employ for selecting $d^{\prime}$ features from the filtered set of variants. We begin by randomly partitioning the variants into buckets of approximately 50 to 100 variants (a hyperparameter) to prevent the loss of information when contexts become too large34. Each bucket is independently passed to the LLM, tasked with selecting the $d^{\prime}$ most relevant variants; we select $d^{\prime}$ at every step in the case that the relevant features are concentrated in a single bucket. The $d^{\prime}$ selected variants from each bucket are merged together, and the process repeats as delineated in Fig. 1. We observe that the selection process is sensitive to the order in which features are presented to the LLM. Thus, for each bucket, we conduct multiple iterations, randomizing the order of variants and using a temperature of 0.3. This approach naturally integrates self-consistency 17 by retaining the top $d^{\prime}$ variants that appear the most across iterations. During the final selection, we enhance the LLM with chain-of-thought (CoT) prompting, increase the temperature to 0.7, and increase the number of iterations to ten.
@@
is the second strategy we employ. Starting with the filtered set of variants, we task the LLM with identifying the single most relevant variant with CoT. After each selection, the chosen variant is removed and the process repeats. Initially, we perform this extraction without any ensembling; the top few features are easy for the LLM to identify, but the task becomes more challenging as it becomes ambiguous which variants are more significant. After selecting a few, we start to apply self-consistency: repeating the extraction several times, in increasing amounts as we near the end of the selection. We find that this also mitigates the LLM’s tendency to return the features that were already selected. In cases where the LLM still fails to identify a valid variant, which occurs frequently towards the end, we implement an exception handling mechanism that retries the selection process.
@@
[…]
Formally, we transform our dataset $D_{S^{\prime}}$ into $K$ transformed datasets $D_{k}=\{(\mathbf{x}^{i}_{S^{\prime}_{k}},y^{i})\}^{N}_{i=1}$, where $S^{\prime}_{k}\supseteq S^{\prime}$. Each transformed dataset is created by prompting the LLM $\mathcal{T}$ with $p^{\prime}$, which includes a serialized representation of selected examples $R=\{(\mathbf{x}^{i}_{S^{\prime}},y^{i})\}_{i\in R\subseteq D_{S^{\prime}}}$, to enhance the LLM with context 31. We define this function $\mathtt{Serialize}$ to convert each row to a textual description (e.g. “The $s_{1}$ variant of the person has $x^{i}_{1}$ minor alleles... The $s_{d^{\prime}}$ variant of the person has $x^{i}_{d^{\prime}}$ minor alleles.”) as LLMs usually prefer natural language. Since LLMs can be sensitive to the input, we examine various serialization templates and prompts. To address gaps in the LLM’s prior knowledge, we also explore retrieval-augmentation to supplement the variant IDs with its associated gene 36.
The LLM $\mathcal{T}$ outputs the new features that are added back onto $S^{\prime}$ to create $S^{\prime}_{k}$:
$S^{\prime}_{k}=S^{\prime}\cup\mathcal{T}\left(S^{\prime},p^{\prime};\theta\right),\text{where }p^{\prime}\text{ includes }\{\mathtt{Serialize}(\mathbf{x}^{i}_{S^{\prime}},y^{i},S^{\prime})\}_{i\in R}$ (2)
We then train $K$ models, =$\mathcal{F}=\{f_{k}\}^{K}_{k=1}$, each on a dataset $D_{k}$, aiming to capture different hypotheses about how the variants relate to the phenotype. During inference, for input $\mathbf{x}$, we average the class probabilities $\mathbf{p}(f_{k}(\mathbf{x}))$ from each model $f_{k}$. The final prediction $\hat{y}$ is made by selecting the class $i$ with the highest averaged probability:
@@
Our aim is to harness the diverse feature representations generated by the LLM, reducing the risk of overfitting in low-shot settings by anchoring each feature set in the knowledge embedded within LLMs rather than the limited data. To uphold interpretability, we limit our feature construction to interaction terms. Notably, our method is model-agnostic, de-coupled from the classifier. We train two models, Random Forest and Logistic Regression, on the transformed datasets to evaluate the quality of the constructed features. We discuss key steps of our method in depth below.
@@
Traditionally, we enforce the output of the LLM to be structured e.g. JSON. However, we deliberately allow the LLM to freely generate its response, recognizing that enforcing a rigid structure can diminish the depth and quality of the LLM’s CoT reasoning 30. To guide the LLM, we provide an example of a “correctly” engineered feature 37, which we self-generate with an LLM and then manually verify.
@@
To mitigate overfitting, we repeat the feature engineering process $K$ times to generate all $D_{k}$ which we’ll use to train $K$ classifiers. Similar to FeatLLM 22, we incorporate additional ensembling and order bias mitigation strategies such as bagging and order shuffling, where we pass a random subset of $|R|\leq N$ samples to the LLM; we limit the number of samples $|R|$ to $16$. We do this to further diversify the LLM output for each iteration, and simultaneously avoid exorbitant usage of the context window. Our free-flow reasoning approach further contributes to the diversity of the output, as we naturally allow the LLM to determine the number of features that will be constructed during its generation. By setting a non-zero temperature of 1, we ensure that each iteration further produces a varied set of features, especially in type. The resulting $k$ transformed datasets form the basis for the ensembled model.
Due to the page limit, we direct readers to our GitHub repo, which contains the detailed prompts and relevant hyperparameters used in the entire pipeline. The source code of FreeForm is provided to promote reproducibility.
Results
@@
Our experiments involve two datasets: the Genomic Ancestry Dataset38 and the Hereditary Hearing Loss Dataset39. The Genomic Ancestry Dataset is derived from the 1000 Genomes Project (1KGP). We focus on determining the superpopulation ancestry phenotype (African, American, East Asian, European, and South Asian). In particular, we used a curated set of 10,000 SNPs predefined by GRAF to pinpoint ancestry markers 40, and we’ll discuss how we addressed the issues that arise from their quality control (QC). After QC and preprocessing, the dataset includes 2,403 subjects and 8,688 variants as columns in the rsID format, a standard identifier used by dbSNP.
The Hereditary Hearing Loss Dataset is considerably smaller, comprising of 1,209 subjects and 144 variants as columns, employing the HGVS nomenclature system. This dataset is notably imbalanced, with approximately 75.9% of the samples classified as “Yes” (indicating the presence of hereditary hearing loss) and 24.1% as “No.” To our best knowledge, these two are the only open-access genotype datasets available online.
@@
For feature selection, we conduct comparisons with four baseline methods. The first three baselines are conventional machine learning approaches: (1) LASSO, (2) PCA and (3) RF-based Gini Importance where we fit a Random Forest on the training data and rank the features by their Gini importance. In our case, the specification of a fixed number of features is required for our study. LASSO, however, reduces an arbitrary number of coefficients to zero during model training. Similarly, PCA provides loadings along principal axes and has been widely used in the genetic field to select variants. To adapt these methods to our requirements, we select the $d^{\prime}$ features with the largest coefficients in the highest-performing LASSO model. For PCA, we take the top $d^{\prime}$ loadings of the first principle component. Lastly, we include (4) LLM-Select20 (we use their LLM-RANK prompts).
For feature engineering, we compared our approach against five baselines. The first three are traditional machine learning methods: (1) Logistic Regression, (2) Random Forest and (3) XGBoost. We also include recent baselines, (4) TabPFN 41, a foundation model for tabular data, and (5) FeatLLM22 which also leverages LLMs to do feature engineering but limits themselves to conditional rules (e.g. variant 1 is $>$ 0) for a linear model.
@@
Our FreeForm framework utilizes GPT-4o (2024-05-13) as the primary LLM backbone, particularly for tasks requiring advanced reasoning capabilities, such as automating feature engineering and selecting relevant features. For more routine tasks, including parsing output and writing Python functions, we employ GPT-3.5-turbo which offers a cost-effective solution that meets the performance requirements for these functions. These models are called upon using the OpenAI API which only requires internet access.
For feature engineering, we employ an ensemble of $K=20$ models, striking a balance between cost-effectiveness and model performance, noting that performance gains diminish beyond this point. In replicating baselines, machine learning models were implemented using Python’s scikit-learn library. Hyperparameters were optimized using grid search and $k$-fold cross-validation, with $k$ set to either 2 or 4, ensuring that the training set includes at least one example of each class. For other methods, such as FeatLLM and TabPFN, we used default parameters with slight adjustments for fair comparison (e.g., using 15 conditions instead of 10 for FeatLLM). Also, for the evaluation of all feature engineering methods, we choose one of the feature sets generated by hierarchical selection.
@@
Figure 3: Evaluation of Feature Selection on Ancestry and Hearing Loss
Our findings indicate that LLM-driven methods significantly outperform data-driven approaches for feature selection in low-shot regimes, achieving gains of up to approximately 20%. Notably, in the genomic ancestry task, LASSO requires 80 shots to achieve similar results to what our framework achieves with just 10 shots. In hearing loss, we observe that the performance gap is smaller and our advantage remains until 16 shots, when using Random Forests. This discrepancy is likely due to the limited presence of variants within the dataset that the LLM has knowledge of.
For feature engineering, we conduct our experiments five times with cross-validation, limiting the evaluation to few-shot settings where $N<120$. In Fig 4, FreeForm consistently consistently ranks at or above baseline models. In the genomic ancestry task, our framework improves the performance of both Logistic Regression and Random Forest, especially in the lower-shot scenarios, and outperforms recent models such as FeatLLM and TabPFN. However, as the number of shots approaches 80, the gap between our methods and the baselines decreases. In the hearing loss task, our framework notably enhances the performance of Logistic Regression on higher-shot scenarios. While Random Forest does not benefit from the engineered features, our approach remains competitive. This outcome suggests limited effectiveness of the interaction terms we engineer for complex models like Random Forests.
Figure 4: Evaluation of Feature Engineering on Ancestry and Hearing Loss